OpenStudy (anonymous):

When trying to solve an equation I'm getting stuck on a step. Rewrite equation in vertex form. y=x^2+4x+3 x^2+4 b/2=4/2=2 x^2+4+(2)^2=x^2+4+4 Now i'm stuck is the next line y=x^2+4x+4-4+3 or y=x^2+4x+4+3-4

OpenStudy (jdoe0001):

\(\bf y=x^2+4x+4+3-4\implies y=(x^2+4x+4)+3-4\\ \quad \\ y=(x^2+4x+2^2)-1\implies y=(x+2)^2-1\)